# Continuous Diffusion for Mixed-Type Tabular Data

### v3

@@ 3.2 Homogenization of Data Types
Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ denote the time-weighted MSE (i.e., score matching) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e., score interpolation) loss of the $j$-th categorical feature. Naturally, the two losses are defined on different scales. This leads to an unintended importance weighting of features in the generative process (Ma et al., 2020). We assume that an unconditional model should a priori, i.e., without having any information, be indifferent between all features. This reflects the state of the model at the terminal timestep $t=1$ in the diffusion process.
Formally, we aim to find calibrated losses, $\mathcal{L}^{\ast}_{\text{MSE}}$ and $\mathcal{L}^{\ast}_{\text{CE}}$ for all continuous features $i$ and categorical features $j$, such that
$\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=\mathbb{E}[\mathcal{L}^{\ast}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]=1.$ (5)
For continuous features, $\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=1$ follows from standardizing $x_{\text{cont}}^{(i)}$ to zero mean and unit variance. For categorical features, we compute the normalization constant $\mathbb{E}[\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]$ directly as the CE of each predicted class in proportion to its empirical distribution in the train set (see Appendix A). We then average the calibrated losses to derive the joint loss function at a given timestep:
$\mathcal{L}(t)=\frac{1}{K}\Bigl{[}\sum_{i=1}^{K_{\text{cont}}}\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)+\sum_{j=1}^{K_{\text{cat}}}\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},t)\Bigr{]},$ (6)
where $K=K_{\text{cont}}+K_{\text{cat}}$.
@@
The loss calibration and the multiple data modalities have implications for the optimal initialization of the score model. We aim to initialize all feature-specific losses at one. We therefore initialize the output layer weights to zero (like in image diffusion models) and the output biases for continuous features to zero, and rely on the timestep weights of the EDM parameterization (Karras et al., 2022) to achieve a unit loss for all $t$. For the categorical features, we initialize the biases to match the category’s empirical probability in the training set (see Appendix B).
@@
The initial equal importance across all timesteps will naturally change over the course of training. We employ a normalization scheme for the average diffusion loss (Karras et al., 2023; Kingma & Gao, 2023) to allow for changes in relative importance among features but ensure equal importance of all timesteps throughout training. To do so, we learn the time-specific normalization term $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$. This ensures a consistent gradient signal and can be implemented by training a neural network to predict $\mathcal{L}(t)$ alongside our diffusion model (for details see Appendix C).
@@ 3.3 Noise Schedules
Since the optimal noise schedule of one feature impacts the noise schedules of other features, and different data types have different sensitivities to additive noise, we introduce feature-specific or type-specific noise schedules. For instance, given the same embedding dimension, more noise is needed to remove the same amount of signal from embeddings of features with fewer classes. Likewise, a delayed noise schedule for one feature might improve sample quality as the model can rely on other correlated features that have been (partially) generated first. We make the noise schedules learnable, and therewith adaptive to avoid the reliance on designs for other data modalities.
@@
According to Equation 1, and following the EDM parameterization (Karras et al., 2022), we define the diffusion process of the $i$-th continuous feature as
$\mathrm{d}x_{{}_{\text{cont}}}^{(i)}=\sqrt{2\Bigl{[}\frac{d}{dt}h_{\text{cont},i}(t)\Bigr{]}h_{\text{cont},i}(t)}\mathrm{d}w_{t}^{(i)},$ (7)
and likewise the trajectory of the $j$-th categorical feature as
$\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}=\sqrt{2\Bigl{[}\frac{d}{dt}h_{\text{cat},j}(t)\Bigr{]}h_{\text{cat},j}(t)}\mathrm{d}{\mathbf{w}}_{t}^{(j)},$ (8)
where ${\mathbf{x}}_{\text{cat}}^{(j)}$ is the $d$-dimensional embedding of $x_{\text{cat}}^{(j)}$ in Euclidean space. The feature-specific noise schedules $h_{\text{cont},i}(t)$ and $h_{\text{cat},j}(t)$ represent the standard deviations of the added Gaussian noise such that $\sigma_{\text{cont},i}(t)=h_{\text{cont},i}(t)$ and $\sigma_{\text{cat},j}(t)=h_{\text{cat},j}(t)$. Thus, each continuous feature and each embedded categorical feature is affected by a distinct noise schedule.
@@
Based on Dieleman et al. (2022), we aim to learn a noise schedule $h_{k}:t\mapsto\sigma$ for all $K=K_{\text{cont}}+K_{\text{cat}}$ features. Note that $t\in[0,1]$, and with pre-specified minimum and maximum noise levels, we can scale $\sigma_{k}$ to lie in $[0,1]$ as well, without loss of generality. We will learn the feature-specific loss given the noise level, $F_{k}:\sigma_{k}\mapsto\ell_{k}$, alongside the score model, with $\ell_{k}$ the relevant (not explicitly weighted) training loss for the $k$-th feature. Then, our mapping of interest is $h_{k}=\tilde{F}_{k}^{-1}$, that is, the normalized and inverted function $F_{k}$. This encourages the relation between $t$ and $\ell_{k}$ to be linear.
Higher noise levels imply a lower signal-to-noise ratio, and therefore a larger incurred loss for the score model. Accordingly, $F_{k}$ must be a monotonically increasing and S-shaped function. We let $F_{k}=\gamma_{k}F_{\mathrm{d.a.log},k}(\sigma_{k})$ where $\gamma_{k}>0$ is a scaling factor that at $t=1$ enables fitting a loss $\ell_{k}>1$ early on in the training process, and a loss $\ell_{k}<1$ in case conditioning information is included. Further, we use the cdf of the domain-adapted Logistic distribution $F_{\mathrm{d.a.log},k}(\sigma_{k})$, where the input is pre-processed via a Logit function, with parameters $10<\mu_{k}<$ 1 (the location of the inflection point) and $\nu_{k}\geq 1$ (the steepness of the curve). Figure 2 illustrates the effect of the location parameter. The implicit importance of the noise levels is conveniently represented by the corresponding pdf $f_{\mathrm{d.a.log},k}$. To normalize and invert $F_{k}$, we set $\gamma_{k}=1$ and and directly utilize the quantile function $F_{\mathrm{d.a.log},k}^{-1}$. The detailed derivation of all relevant functions is given in Appendix D.
Our functional choice has several advantages. First, each noise schedule can be evaluated exactly without the need for approximations and only requires three parameters. Second, these parameters are well interpretable in the diffusion context and provide information on the inner workings of the model. For instance, for $\mu_{1}<\mu_{2}$, the model starts generating feature 2 before feature 1 in the reverse process. Third, the proposed functional form is less flexible than the original piece-wise linear function (Dieleman et al., 2022) such that an exponential moving average on the parameters is not necessary, and the fit is more robust to “outliers” encountered during training.
We use the adaptive noise schedules during both training and generation. We derive importance weights from $f_{\mathrm{d.a.log},k}$ to fit $h_{k}$ to avoid biasing the noise schedule to timesteps that are frequently sampled during training. Type-specific noise schedules refer to learning two functions $F_{1}$ and $F_{2}$ that predict the respective average loss over all features of a data type. Examples of learned noise schedules are given in Appendix O.
@@ 3.4 Additional Customization to Tabular Data
In the diffusion process, we add noise directly to the continuous features but to the embeddings of categorical features. We generally need more noise to remove all signal from the categorical representations. We therefore define type-specific minimum and maximum noise levels: For categorical features, we let $\sigma_{\text{cat},\min}=0.1$ and $\sigma_{\text{cat},\max}=100$; for continuous features, we set $\sigma_{\text{cont},\min}=0.002$ and $\sigma_{\text{cont},\max}=80$ (see Karras et al., 2022).
Lastly, an uninformative initialization of the adaptive noise schedules requires to set $\mu_{k}=0.5,\nu_{k}\approx 1$ and $\gamma_{k}=1$ such that $F_{\mathrm{d.a.log},k}$ corresponds approximately to the cdf of a uniform distribution. We can improve this with a more informative prior: In the image domain, diffusion models allocate substantial capacity towards generating the high level structure before generating details at lower noise levels. In tabular data, the location of features in the data matrix, and therefore the high level structure, is fixed. Instead, we are interested in generating details as accurately as possible, as these influence, for instance, subtle correlations among features. Note that the inflection point, $\mu_{k}$, of our adaptive noise schedule corresponds to the proportion of high (normalized) noise levels (i.e., $\sigma_{k}\geq 0.5$) in the distribution. Therefore, we adjust the initial noise schedules such that low noise levels ($\sigma_{k}<0.5$) are weighted by a factor of 3 relative to high noise levels ($\sigma_{k}\geq 0.5$) (see Figure 2). The proportion of high noise levels is decreased to $\mu_{k}=1/4$. We let $\nu_{k}\approx 1$ for a dispersed initial probability mass and initialize the scaling factor to $\gamma_{k}=1$.
@@ 4 Experiments
We use a diverse benchmark set of state-of-the-art generative models for mixed-type tabular data. This includes SMOTE (Chawla et al., 2002), ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), CoDi (Lee et al., 2023), TabSyn (Zhang et al., 2024). Each model follows a different design and/or modeling philosophy. Note that CoDi is an extension of STaSy (Kim et al., 2023, the same group of authors) that has shown to be superior in performance. For scaling reasons, ForestDiffusion (Jolicoeur-Martineau et al., 2024) is not an applicable benchmark. [^1] Jolicoeur-Martineau et al. (2024) report in the appendix that they used 10-20 CPUs with 64-256 GB of memory for datasets with a median number of 540 observations. With the suggested hyperparameters (for improved efficiency) and 64 CPUs, the model took approx. 500 min of training on the relatively small nmes data. Note that the model estimates $KT$ separate models, with $K$ being the number of features and $T$ the noise levels. Therefore, we consider ForestDiffusion to be prohibitively expensive for higher-dimensional data generation. Further details on the respective benchmark models and their implementations are provided in Appendix F and Appendix G. We provide an in-depth comparison of CDTD to the diffusion-based baselines in Appendix N. To keep the comparison fair, we use the same architecture for CDTD as TabDDPM (the latter has also been adopted by TabSyn), with minor changes to accommodate the different inputs (see Appendix J).
@@
We systematically investigate our model on eleven publicly available datasets. The datasets vary in size, prediction task (regression vs. binary classification [^2] For ease of presentation, we only analyze binary targets. However, CDTD trivially extends to targets with multiple classes.), number of continuous and categorical features and their distributions. The number of categories for categorical features varies significantly across datasets (for more details, see Appendix E). We remove observations with missings in the target or any of the continuous features and encode missings in the categorical features as a separate category. All datasets are split in train (60%), validation (20%) and test (20%) partitions, hereinafter denoted $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, respectively. For classification tasks, we use stratification with respect to the outcome, we condition the model on the outcome during training and generation, and use the train set proportions for generation. In a last post-processing step, we round the integer-valued continuous features after generation for all models.
@@ 4.1 Evaluation Metrics
In our experiments, we follow conventions from previous papers and use four sample quality criteria, which we assess using a comprehensive set of measures. All metrics are averaged over five random seeds that affect the generative process, which samples synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$.
@@
We follow the conventional train-synthetic-test-real strategy (see, Borisov et al., 2023; Liu et al., 2023; Kotelnikov et al., 2023; Kim et al., 2023; Xu et al., 2019; Watson et al., 2023). Hence, we train a group of models, consisting of a (logistic/ridge) regression, a random forest and a catboost model, on the data-specific prediction task (the corresponding hyperparameter settings are reported in Appendix I). We compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance when trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$. We subsample $\mathcal{D}_{\text{train}}$ in case of more than 50 000 observations to upper-bound the computational load. The results are averaged over ten different model seeds (in addition to the five random seeds that impact the sampling process). For regression tasks, we consider the RMSE and for classification tasks, the macro-averaged F1 and AUC scores. We only report $\lvert\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})-\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\rvert$ in the main part of this paper. An absolute difference close to zero, that is, synthetic and real data induce the same performance, indicates that the generative model performs well.
@@
For each generative model, we report the accuracy of a catboost model that is trained to distinguish between real and generated (fake) samples (Borisov et al., 2023; Liu et al., 2023; Zhang et al., 2024). First, we subsample the real data subsets, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, to a maximum of 25 000 data samples to limit evaluation time. Then, we construct $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ with equal proportions of real and fake samples. We tune each catboost model on $\mathcal{D}_{\text{valid}}^{\text{detect}}$ and report the accuracy of the best-fitting model on $\mathcal{D}_{\text{test}}^{\text{detect}}$ (see Appendix H for details). A (perfect) detection score of 0.5 indicates the model is unable distinguish fake from real samples.
@@
We aim to assess the statistical similarity between real and generated data at both the feature and sample levels. We largely follow Zhao et al. (2021) and compare: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance between pair-wise correlation matrices. We use the Pearson correlation coefficient for two continuous features, the Theil uncertainty coefficient for two categorical features, and the correlation ratio for mixed types. Similar metrics for the evaluation of statistical similarity have been used by Zhang et al. (2024).
Distance to closest record.
That is, the minimum Euclidean distance of a generated data point to any observation in $\mathcal{D}_{\text{train}}$ (Borisov et al., 2023; Zhao et al., 2021). We one-hot encode categorical features and standardize all features to zero mean and unit variance to ensure each feature contributes equally to the distance. We compute the average distance to closest record (DCR) as a robust estimate. For brevity, we report the absolute difference of the DCR of the synthetic data and the DCR of the real test set. A good DCR value, indicating both realistic and sufficiently private data, should be close to zero.

### v4

@@ 3.2 Homogenization of Data Types
Let $\mathcal{L}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)$ be the time-weighted MSE (i.e., score matching) loss of the $i$-th continuous feature at a single timestep $t$, and $\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},t)$ the CE (i.e., score interpolation) loss of the $j$-th categorical feature. Naturally, the two losses are defined on different scales. This leads to an unintended importance weighting of features in the generative process (Ma et al., 2020). To solve this, we observe that an unconditional generative model should a priori, i.e., without having any information, be indifferent between all features. For diffusion models this reflects the state of the model at the terminal timestep $t=1$.
Definition 1 (Calibrated losses).
Scaled losses $\mathcal{L}^{\ast}_{\text{MSE}}$ and $\mathcal{L}^{\ast}_{\text{CE}}$ are called calibrated if
$\mathbb{E}[\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},1)]=\mathbb{E}[\mathcal{L}^{\ast}_{\text{CE}}(x_{\text{cat}}^{(j)},1)]=1,$
for all continuous features $i$ and categorical features $j$.
Assumption 1 (No information at $t=1$).
The noise level $\sigma(1)$ is high enough such that the input to the score model cannot be distinguished from noise, i.e., the model has no information.
Proposition 1 (Homogenization of feature-specific losses).
Under Assumption 1, if each $x_{\text{cont}}^{(i)}$ has unit variance and $Z_{j}$ is the feature-specific entropy of categorical feature $j$, then the losses $\mathcal{L}^{\ast}_{\text{MSE}}=\mathcal{L}_{\text{MSE}}$ and $\mathcal{L}^{\ast}_{\text{CE}}=\mathcal{L}_{\text{CE}}(x_{\text{cat}}^{(j)},1)/Z_{j}$ are calibrated for all continuous features $i$ and categorical features $j$. See Appendix B for a detailed proof.
Given Assumption 1 and Proposition 1, we can derive a joint loss function without unintended importance weighting by averaging the $K=K_{\text{cont}}+K_{\text{cat}}$ calibrated losses at a given $t$:
$\mathcal{L}(t)=\frac{1}{K}\Bigl{[}\sum_{i=1}^{K_{\text{cont}}}\mathcal{L}^{\ast}_{\text{MSE}}(x_{\text{cont}}^{(i)},t)+\sum_{j=1}^{K_{\text{cat}}}\mathcal{L}_{\text{CE}}^{\ast}(x_{\text{cat}}^{(j)},t)\Bigr{]}.$ (5)
Implications for the score model initialization.
@@
The loss calibration and the multi-modality of the data have implications for the optimal initialization of the score model. To match the loss calibration, we aim to initialize the model such that all feature-specific losses are one. We therefore initialize the output layer weights and biases for continuous features to zero and rely on the timestep weights of the EDM parameterization (Karras et al., 2022) to achieve a unit loss for all $t$. For the categorical features, we initialize the biases to match each category’s empirical log probability in the training set (see Appendix C for details).
Weighting across time.
@@
The initial equal importance across all $t$ will change over the course of training. To allow for changes in the relative importance among features but ensure equal importance of all timesteps throughout training, we employ a normalization scheme for the average diffusion loss (Karras et al., 2024; Kingma & Gao, 2023). Specifically, we learn the time-dependent normalization $Z(t)$ such that $\mathcal{L}(t)/Z(t)\approx 1$. This ensures a consistent gradient signal and can be implemented by training a neural network to predict $\mathcal{L}(t)$ alongside our diffusion model (see Appendix D for details).
@@ 3.3 Noise Schedules
Evidently, the noise schedule of one feature impacts the optimality of noise schedules for other features, and different data types have different sensitivities to additive noise. For instance, given the same embedding dimension, more noise may be needed to remove the same amount of signal from embeddings of features with fewer classes (see Appendix A). Likewise, a delayed noise schedule for one feature might improve sample quality as the model can rely on other correlated features that have been (partially) generated first. Therefore, we introduce feature-specific or type-specific noise schedules. We make the noise schedules learnable, and therewith adaptive to avoid the reliance on designs for other data modalities.
@@
Following Equation 1, we define the diffusion process of $x_{{}_{\text{cont}}}^{(i)}$ as
$\mathrm{d}x_{{}_{\text{cont}}}^{(i)}=\sqrt{2\Bigl{[}\frac{\mathrm{d}}{\mathrm{d}t}h_{\text{cont},i}(t)\Bigr{]}h_{\text{cont},i}(t)}\mathrm{d}w_{t}^{(i)},$ (6)
and likewise the trajectory of ${\mathbf{x}}_{\text{cat}}^{(j)}$ as
$\mathrm{d}{\mathbf{x}}_{\text{cat}}^{(j)}=\sqrt{2\Bigl{[}\frac{\mathrm{d}}{\mathrm{d}t}h_{\text{cat},j}(t)\Bigr{]}h_{\text{cat},j}(t)}\mathrm{d}{\mathbf{w}}_{t}^{(j)},$ (7)
where ${\mathbf{x}}_{\text{cat}}^{(j)}\in{\mathbb{R}}^{d}$ is the embedding of $x_{\text{cat}}^{(j)}$ in Euclidean space. The noise schedules $h_{\text{cont},i}(t)$ and $h_{\text{cat},j}(t)$ represent the feature-specific standard deviations $\sigma_{\text{cont},i}(t)$ and $\sigma_{\text{cat},j}(t)$ of the added Gaussian noise, respectively. Thus, each feature is affected by a distinct adaptive noise schedule. On the other hand, type-specific noise schedules involve only two functions, $h_{\text{cat}}(t)$ and $h_{\text{cont}}(t)$, such that all features of the same type are affected by the same noise schedule.
@@
We aim to learn a noise schedule $h_{k}:t\mapsto\sigma_{k}$ for each feature $k$. Inspired by Dieleman et al. (2022), we learn a function $F_{k}$ that predicts the feature-specific (not explicitly weighted) loss $\ell_{k}$ given the noise level $\sigma_{k}$. By normalizing and inverting $F_{k}$, we achieve the mapping of interest $h_{k}=\tilde{F}_{k}^{-1}$. This encourages the relation between $t$ and $\ell_{k}$ to be linear.
Higher noise levels imply a lower signal-to-noise ratio and therefore a larger incurred loss for the score model. Accordingly, $F_{k}$ must be a monotonically increasing and S-shaped function. We let $F_{k}=\gamma_{k}F_{\mathrm{d.a.log},k}(\sigma_{k})$ where $\gamma_{k}>0$ is a scaling factor that enables fitting a loss $\ell_{k}>1$ at $t=1$ near the start of training, and a loss $\ell_{k}<1$ in case conditioning information is included. Further, we use the cdf of the domain-adapted Logistic distribution $F_{\mathrm{d.a.log},k}(\sigma_{k})$, where the input is pre-processed via a Logit function, with parameters $10<\mu_{k}<$ 1 (the location of the inflection point) and $\nu_{k}\geq 1$ (the steepness of the curve). Note that with pre-specified minimum and maximum noise levels, we can always scale $\sigma_{k}$ to lie in $[0,1]$. Figure 2 illustrates the effect of the location parameter. The implicit importance of the noise levels is conveniently represented by the pdf $f_{\mathrm{d.a.log},k}$. To normalize and invert $F_{k}$, we set $\gamma_{k}=1$ and directly use the closed-form quantile function $F_{\mathrm{d.a.log},k}^{-1}$. The detailed derivation of all relevant functions is given in Appendix E. To avoid biasing the noise schedule to frequently sampled timesteps during training, we derive importance weights from $f_{\mathrm{d.a.log},k}$ when fitting $h_{k}$. We use the adaptive noise schedules during both training and generation, and give examples of learned noise schedules in Appendix P.
Our functional choice has several advantages. First, each noise schedule can be evaluated exactly without the need for approximations and only requires three parameters. Second, these parameters are well interpretable in the diffusion context and provide information on the inner workings of the model. For instance, for $\mu_{1}<\mu_{2}$, the model starts generating feature 2 before feature 1 in the reverse process. Third, the proposed functional form is less flexible than the original piece-wise linear function (Dieleman et al., 2022) such that an EMA on the parameters is not necessary, and the fit is more robust to “outliers” encountered during training. This is crucial when using a feature-specific specification as opposed to a single noise schedule.
@@ 3.4 Additional Customization to Tabular Data
In the diffusion process, we add noise directly to the continuous features but to the embeddings of categorical features. We generally need more noise to remove all the signal from the categorical representations. We therefore define type-specific minimum and maximum noise levels: For categorical features, we let $\sigma_{\text{cat},\min}=0$ and $\sigma_{\text{cat},\max}=100$; for continuous features, we set $\sigma_{\text{cont},\min}=0$ and $\sigma_{\text{cont},\max}=80$.
Lastly, an uninformative initialization of the adaptive noise schedules requires $\mu_{k}=0.5,\nu_{k}\approx 1$ and $\gamma_{k}=1$ such that $F_{\mathrm{d.a.log},k}$ corresponds approximately to the cdf of a uniform distribution. We can improve upon this with a more informative prior: As opposed to images, in tabular data the location of features in the data matrix, and therefore the high-level structure, is fixed. Instead, we are interested in generating details as accurately as possible. Note that the inflection point, $\mu_{k}$, corresponds to the proportion of high noise levels (i.e., $\sigma_{k}\geq 0.5\cdot(\sigma_{\max}+\sigma_{\min})$) in the distribution. Therefore, we empirically choose $\mu_{k}=1/4$ such that low noise levels (i.e., $\sigma_{k}<0.5\cdot(\sigma_{\max}+\sigma_{\min})$) are initially presumed to be three times more important than high noise levels (see Figure 2). We let $\nu_{k}\approx 1$ for a dispersed initial probability mass and initialize the scaling factor to $\gamma_{k}=1$.
@@ 4 Experiments
We use a diverse benchmark set of state-of-the-art generative models for mixed-type tabular data. This includes SMOTE (Chawla et al., 2002), ARF (Watson et al., 2023), CTGAN (Xu et al., 2019), TVAE (Xu et al., 2019), TabDDPM (Kotelnikov et al., 2023), CoDi (Lee et al., 2023) and TabSyn (Zhang et al., 2024). Each model follows a different design and/or modeling philosophy. Note that CoDi is an extension of STaSy (Kim et al., 2023, the same group of authors) that has shown to be superior in performance. For scaling reasons, ForestDiffusion (Jolicoeur-Martineau et al., 2024) is not an applicable benchmark. [^1] Jolicoeur-Martineau et al. (2024) report that they used 10-20 CPUs with 64-256 GB of memory for datasets with a median number of 540 observations. With the suggested hyperparameters (for improved efficiency) and 64 CPUs, the model took approx. 500 min of training on the small nmes data. Note that the model estimates $KT$ separate models, with $K$ being the number of features and $T$ the noise levels. Therefore, we consider ForestDiffusion to be prohibitively expensive for higher-dimensional data generation. Further details on the respective benchmark models and their implementations are provided in Appendix G and Appendix H, respectively. We provide an in-depth comparison of CDTD to the diffusion-based baselines in Appendix O. To keep the comparison fair, we use the same internal architecture for CDTD as TabDDPM (which was also adopted by TabSyn), with minor changes to accommodate the different inputs (see Appendix K).
@@
We systematically investigate our model on 10 publicly available datasets (see Appendix F for details). The datasets vary in size, prediction task (regression vs. binary classification [^2] For ease of presentation, we only analyze binary targets. However, CDTD trivially extends to targets with multiple classes.), number of continuous and categorical features and their distributions. The number of categories for categorical features varies significantly across datasets. We remove observations with missings in the target or any of the continuous features and encode missings in the categorical features as a separate category. All datasets are split in train (60%), validation (20%) and test (20%) partitions, hereinafter denoted $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, respectively. For classification tasks, we use stratification with respect to the outcome. We round the integer-valued continuous features after generation.
@@ 4.1 Evaluation Metrics
In our experiments, we follow conventions from previous papers and use four sample quality criteria, which we assess using a comprehensive set of measures. All metrics are averaged over five random seeds that affect the sampling process of synthetic data $\mathcal{D}_{\text{gen}}$ of size $\min(\lvert\mathcal{D}_{\text{train}}\rvert,50\,000)$.
@@
We follow the conventional train-synthetic-test-real strategy (see, Borisov et al., 2023; Liu et al., 2023; Kotelnikov et al., 2023; Kim et al., 2023; Xu et al., 2019; Watson et al., 2023). We train a logistic/ridge regression, a random forest and a catboost model, on the data-specific prediction task (see Appendix J for details) and then compare the model-averaged real test performance, $\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})$, to the performance when trained on the synthetic data, $\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})$. The results are averaged over ten different model seeds (in addition to the five seeds for the sampling). For regression tasks, we consider the RMSE and for classification tasks, the macro-averaged F1 and AUC scores. We report $\lvert\text{Perf}(\mathcal{D}_{\text{gen}},\mathcal{D}_{\text{test}})-\text{Perf}(\mathcal{D}_{\text{train}},\mathcal{D}_{\text{test}})\rvert$. An absolute difference close to zero is preferable since then synthetic and real data induce the same performance.
@@
For each generative model, we report the accuracy of a catboost model that is trained to distinguish between real and generated (fake) samples (Borisov et al., 2023; Liu et al., 2023; Zhang et al., 2024). First, we subsample the real data subsets, $\mathcal{D}_{\text{train}},\mathcal{D}_{\text{valid}}$ and $\mathcal{D}_{\text{test}}$, to a maximum of 25 000 data samples to limit evaluation time. Then, we construct $\mathcal{D}_{\text{train}}^{\text{detect}},\mathcal{D}_{\text{valid}}^{\text{detect}}$ and $\mathcal{D}_{\text{test}}^{\text{detect}}$ with equal proportions of real and fake samples. We tune each catboost model on $\mathcal{D}_{\text{valid}}^{\text{detect}}$ and report the accuracy of the best-fitting model on $\mathcal{D}_{\text{test}}^{\text{detect}}$ (see Appendix I for details). A (perfect) detection score of 0.5 indicates that the model is unable to distinguish fake from real samples.
@@
Similar to Zhang et al. (2024), we assess the statistical similarity between real and generated data at both the feature and sample levels. We largely follow Zhao et al. (2021) and compare: (1) the Jensen-Shannon divergence (JSD; Lin, 1991) to quantify the difference in categorical distributions, (2) the Wasserstein distance (WD; Ramdas et al., 2017) to quantify the difference in continuous distributions, and (3) the L2 distance between pairwise correlation matrices. We use the Pearson correlation coefficient for two continuous features, the Theil uncertainty coefficient for two categorical features, and the correlation ratio for mixed types.
Privacy.
We compute the distance to closest record (DCR) as the minimum Euclidean distance of a generated data point to any observation in $\mathcal{D}_{\text{train}}$ (Borisov et al., 2023; Zhao et al., 2021). We one-hot encode categorical features and standardize all features to zero mean and unit variance to ensure each feature contributes equally to the distance. We compute the average DCR as a robust estimate. For brevity, we report the absolute difference of the DCR of the synthetic data and the DCR of the real test set. A good DCR value, indicating both realistic and sufficiently private data, should be close to zero.